A stone weighting 0.5kg falls from a height and strikes the ground in 4 seconds. It buries deep in sand in 0.09 seconds. The depth of penetration is (g=10m/s​2)
a) 100m
b) 2.0m
c) 1.8m
d) 0.9m

Grade:12th pass

1 Answers

Eshan
4 years ago
Dear student,

Velocity of stone when it reaches the ground= u+at=40m/s

The stone stops in 0.09seconds. Therefore from third equation of motion,

a=\dfrac{-u}{t}=-\dfrac{40}{0.09}m/s^2
Therefore the distance travelled=-\dfrac{u^2}{2a}=\dfrac{40\times 0.09m}{2}=1.8m